# Five Years of SciCap: What We Learned and Future Directions for Scientific Figure Captioning

### v1

@@ Abstract
The story of SciCap began in late 2019, when we at Penn State started discussing whether we could build a large-scale dataset of real-world scientific figures and captions to support multimodal modeling for scientific visuals. The spark came when Lee walked into Kenneth’s office and asked, “Can we build a CaptionBERT?” Lee knew Kenneth had worked on vision-and-language datasets and models before, including the Visual Storytelling (VIST) dataset, and around that time domain-specific BERT variants, such as SciBERT (beltagy-etal-2019-scibert) and BioBERT (10.1093/bioinformatics/btz682), were showing clear advantages. Their success made us wonder whether scientific figure captions might also benefit from domain-specific modeling, if only we could gather enough real data.
We soon drafted a seed-grant proposal to Penn State outlining two motivations: First, figure captions in scientific papers were often poorly written, frequently lacking the detail or clarity needed for readers to understand a figure without returning to the full text. This seemed like an ideal place for automated assistance. Second, if we wanted to build technology to improve caption quality, there was a fundamental data gap: no large-scale, real-world dataset of scientific figures paired with their captions existed. Most prior datasets were either synthetic or too small to support reliable captioning models (Kafle_2018_CVPR; kahou2017figureqa; 10.1145/3341162.3345601; Chen_2020_WACV). We aimed to fix both problems at once.
The proposal was funded. By 2021, we had constructed and released the first version of the SciCap dataset (hsu-etal-2021-scicap-generating), focusing on graph plots from computer-science and machine-learning arXiv papers. We gathered 295,028 papers published between 2010 and 2020, extracted over 2.17 million figures, identified 416,804 graph plots (19.2%), and after filtering out compound figures, produced a clean set of 133,543 single-panel figures.
In the 2021 SciCap release paper (hsu-etal-2021-scicap-generating), we also evaluated several baseline vision-to-language models. Their performance was strikingly poor: captions were often low-quality, inaccurate, or even outright gibberish. The results confirmed our intuition: scientific figure captioning was far from a solved problem, and real progress would require both better data and better models.
@@ A Highly Contextual Task: Captioning as Contextual Summary, Not Just Visual Description
Soon after we released the first version of SciCap, we learned that Adobe Research had also been exploring scientific figure captioning (10.1145/3442381.3449923). We reached out, and after a few very encouraging conversations, our two groups decided to join forces in Fall 2021. That collaboration ended up shaping the direction of SciCap for years.
@@
Gradually, we noticed a pattern: many scientific figure captions were essentially paraphrases or compressed summaries of the paper’s own descriptions. That observation led to a simple but surprising experiment. We treated figure captioning as a text-summarization problem and fine-tuned a summarization model, Pegasus (zhang2020pegasus), using only the figure-mentioning paragraphs—no images at all. The text-only approach worked remarkably well. We also found that roughly 75% of caption words could be traced directly to those paragraphs or the figure’s OCR.
This work resulted in our INLG 2023 paper, which received the Best Long Paper award and was selected as one of the three finalists for the Best Evaluation Paper award (huang-etal-2023-summaries). It demonstrated that scientific figure captioning behaves much more like contextual summarization than visual description. Furthermore, a small human evaluation in the paper confirmed our motivating observation: many author-written captions were mediocre at best. PhD students rated over 50% of author-written captions as unhelpful. This result highlights the limitations of caption authoring practices and the potential for LLM assistance to improve caption utility for readers.
@@
Against this backdrop, as well as with ChatGPT reshaping the field, we began experimenting with LLMs as automatic judges in early 2023. PhD students and trained undergraduates rated captions generated by different approaches, and we compared their ratings to those produced by LLMs. The results, published at EMNLP 2023 (hsu-etal-2023-gpt), were interesting: LLM judgments aligned well with PhD students’ ratings. Undergraduate ratings, however, diverged. Less-experienced readers cared more about clear take-away messages, while experts prioritized numerical precision and technical detail.
Our study landed at the early wave of the “LLM-as-judge” trend and drew more attention than we expected. Interestingly, a later 2024 study would report that LLM judges sometimes favor their own generations (10.5555/3737916.3740113). Our work avoided that bias by deliberately excluding LLM-generated captions from the evaluation pool.
@@
We launched the first SciCap Challenge in 2023 (Figure 1) at the ICCV workshop Closing the Loop Between Vision and Language. We invited teams around the world to generate captions for an expanded version of our dataset, i.e., the SciCap Challenge dataset, which contained 476,389 single-panel scientific figures across 8 scholarly domains and 5 figure types. [^1] The SciCap Challenge dataset (hsu2025large) included figures from arXiv papers across computer science, physics, math, economics, statistics, quantitative biology, quantitative finance, and electrical engineering. Figures were categorized into node diagrams, equations, graph plots, scatterplots, and bar charts. We limited the per-paper sampling to encourage participation from teams with modest computational budgets. This dataset was more than three times the size of our original 2021 SciCap release. Each figure came with its image, OCR text, figure type, and the paragraphs in the arXiv that explicitly mentioned it. When we designed the specification of the SciCap Challenge, we intentionally made the figure-mentioning paragraphs part of the standard input that every team received. We wanted this design choice to send a clear signal: scientific figure captioning is a highly context-dependent generation task; a good caption cannot be written from an image alone.
Built on the outputs submitted to the 2023 Challenge, we conducted a large-scale human and automatic evaluation study, later published at TACL and presented at EMNLP 2025 (hsu2025large). In this paper, we found that multimodal LLMs (MLLMs), particularly GPT-4V, consistently produced captions preferred by professional editors over all other approaches, including human-written captions. This was true across domains, across figure types, and even for papers published after the models’ claimed training cut-off date. In cases where authors wrote captions carelessly, MLLM-generated captions were often clearly better—a humbling result, but also an exciting one for the potential of assistive tools.
Since 2023, we have continued to run the SciCap Challenge annually—at IJCAI 2024 and at the LM4Sci Workshop at COLM 2025. Over time, this effort has helped form a small yet active community working on scientific figure captioning. Some teams have even returned to participate again after their first participation; we even collaborated with a team to publish their approach (10.1007/978-981-96-8912-5_6). Our very first SciCap paper (hsu-etal-2021-scicap-generating), which originally was a Findings paper without fancy modeling, has been cited over 130 times.
@@
Each year we ran the SciCap Challenge, we faced the same dilemma: automatic evaluation is cheap and fast but unreliable, while human evaluation is reliable but extremely slow and expensive. Across the 2023-2025 challenges, we experimented with different ways to balance this trade-off. In 2023, we gave two awards: one based on ROUGE-2-Normalized scores on the full hidden test set, and another on a curated subset of higher-quality captions, motivated by our finding that models performing well overall could still fail on well-written captions (huang-etal-2023-summaries). However, rankings barely changed across the two sets. In 2024, we switched to human evaluation on a small sample of 200 figures and split the challenge into two tracks: a short-caption track, where systems had to produce captions no longer than the originals for at least 30% of samples, and a long-caption track with the opposite constraint. This time, the results diverged: teams outperformed the original authors of captions in the long-caption track, but authors still performed better in the short-caption track. This could be due to humans’ known preference for longer captions. In 2025, to mitigate this bias, we went with two distinct tracks, one judged by humans and one judged automatically, using the new LaMP-Cap dataset (ng2025lamp) focused on personalized captioning. This time, the results diverged again across two sets. Teams were able to mimic the author’s style per automatic evaluations, but this often hurt caption quality perceived by humans. In fact, human evaluators consistently preferred GPT-4o’s vanilla captions over personalized ones.
@@
Our first step was building SciCapenter, an interactive caption-writing assistant that displays the figure, its figure-mentioning paragraphs, and several AI-generated draft captions, and allows writers to iteratively revise them with quality ratings and contextual guidance (10.1145/3613905.3650738). A pilot study suggested that these suggestions reduced cognitive load and helped writers produce higher-quality text. We then conducted a separate, deeper study in which 18 researchers rewrote captions from their own recently published papers while using multiple LLM-generated captions as starting points (ngunderstanding). Through video-based interaction analysis, we found that writers typically began by copying or adapting AI-generated captions, preferred longer and detail-rich drafts, and relied heavily on the textual–visual alignment of the suggestions. However, they struggled when figures were conceptually complex or when AI suggestions conflicted with disciplinary norms. We also saw broader themes (ng-etal-2025-understanding): caption writing is deeply multimodal, writers often lack linguistic confidence but trust their domain expertise, they benefit from multiple diverse AI suggestions, and caption norms vary dramatically across fields.
@@
In our caption-writing studies (ngunderstanding; ng-etal-2025-understanding), we noticed a recurring gap: even when AI-generated captions were of fine quality, authors often felt they were too generic and did not match their writing style or the stylistic conventions of their target venues. This pushed us to think seriously about personalization. Although personalized text generation is now common with LLMs (zhang2024personalization), the key practical question for our task was: what information can we use to characterize the style of a group of co-authors? Using all papers written by each co-author was an obvious answer, but it was expensive and conceptually messy, since scholarly writing is collaborative and we rarely know which author actually wrote which caption. So we turned to something more grounded and reliable: information within the same paper.
This led us to create LaMP-Cap (ng2025lamp), a dataset derived from the SciCap Challenge dataset and designed specifically for personalized scientific figure captioning using multimodal profiles. For each target figure, LAMP-CAP provides the figure image and its figure-mentioning paragraphs as model input, along with up to three additional figures from the same paper (each with its own image, caption, and paragraphs) as a style profile. Across 110,828 target figures, we tested four LLMs and consistently found that adding these multimodal profiles helped models generate captions closer to the original author-written ones. Human evaluation further confirmed the usefulness of personalization. Interestingly, human judges preferred captions generated with one profile figure over both no-profile and all-profile conditions. This echoes a common dilemma in personalization (zhang2024personalization): a little tailoring improves quality and feels helpful to users, but pushing too hard to mimic the user’s style can hurt the generation quality.
@@
Real-world writing rarely unfolds in neat, fully drafted documents. To help authors, captioning systems must handle partial or missing context for figures in unfinished papers or for standalone visuals circulating on social media, in news, or in advertisements. 18.81% of figures in the original SciCap dataset lacked any identified mentions (huang-etal-2023-summaries). Generating meaningful captions when only the image is available remains one of the field’s most practical yet unsolved problems.
@@
SciCap’s evaluations showed that reader expertise shapes caption preference (hsu-etal-2023-gpt). Undergraduates valued clear takeaways; PhD students cared about methodological precision and numbers. Future models might adapt captions for different audiences, such as experts, students, and practitioners, which will require advances in audience modeling, controllable generation, and implicit preference inference. Building audience-aware captioning systems without relying on extensive user profiling remains an open challenge in personalization.
@@
The experimental results from both the LaMP-CAP dataset (ng2025lamp) and the SciCap Challenge 2025 showed a clear trade-off: strong stylistic imitation can reduce factual accuracy or clarity. This tension between stylistic faithfulness and communicative quality raises important questions about authorship, identity, and the design of human-AI co-writing systems. It also brings up a broader question that some NLP researchers have asked: to support writing, do we actually need personalized text generation, or is generic generation sufficient for most authors to build on?
@@
Most caption research evaluates outputs from the reader’s perspective, not the writer’s. Yet to understand how AI truly helps scientists, we must study caption writing in realistic conditions. Recruiting authors actively drafting their own papers is difficult; proxy studies, e.g., rewriting old captions or captioning others’ figures (ngunderstanding; ng-etal-2025-understanding), only approximate the real task. The Sloan-funded deployment of SciCapenter opens a rare opportunity to observe genuine writing-in-progress (Huang2024_ScientificCaptions), but conducting such ecologically valid studies remains expensive and methodologically hard.

### v2

@@ Abstract
The story of SciCap began in late 2019, when we at Penn State started discussing whether we could build a large-scale dataset of real-world scientific figures and captions to support multimodal modeling for scientific visuals. The spark came when Lee walked into Kenneth’s office and asked, “Can we build a CaptionBERT?” Lee knew Kenneth had worked on vision-and-language datasets and models before, including the Visual Storytelling (VIST) dataset, and around that time domain-specific BERT variants, such as SciBERT (Beltagy et al. 2019) and BioBERT (Lee et al. 2019), were showing clear advantages. Their success made us wonder whether scientific figure captions might also benefit from domain-specific modeling, if only we could gather enough real data.
We soon drafted a seed-grant proposal to Penn State outlining two motivations: First, figure captions in scientific papers were often poorly written, frequently lacking the detail or clarity needed for readers to understand a figure without returning to the full text. This seemed like an ideal place for automated assistance. Second, if we wanted to build technology to improve caption quality, there was a fundamental data gap: no large-scale, real-world dataset of scientific figures paired with their captions existed. Most prior datasets were either synthetic or too small to support reliable captioning models (Kafle et al. 2018; Kahou et al. 2017; Chen et al. 2019, 2020). We aimed to fix both problems at once.
The proposal was funded. By 2021, we had constructed and released the first version of the SciCap dataset (Hsu et al. 2021), focusing on graph plots from computer-science and machine-learning arXiv papers. We gathered 295,028 papers published between 2010 and 2020, extracted over 2.17 million figures, identified 416,804 graph plots (19.2%), and after filtering out compound figures, produced a clean set of 133,543 single-panel figures.
In the 2021 SciCap release paper (Hsu et al. 2021), we also evaluated several baseline vision-to-language models. Their performance was strikingly poor: captions were often low-quality, inaccurate, or even outright gibberish. The results confirmed our intuition: scientific figure captioning was far from a solved problem, and real progress would require both better data and better models.
@@ A Highly Contextual Task: Captioning as Contextual Summary, Not Just Visual Description
Soon after we released the first version of SciCap, we learned that Adobe Research had also been exploring scientific figure captioning (Qian et al. 2021). We reached out, and after a few very encouraging conversations, our two groups decided to join forces in Fall 2021. That collaboration ended up shaping the direction of SciCap for years.
@@
Gradually, we noticed a pattern: many scientific figure captions were essentially paraphrases or compressed summaries of the paper’s own descriptions. That observation led to a simple but surprising experiment. We treated figure captioning as a text-summarization problem and fine-tuned a summarization model, Pegasus (Zhang et al. 2020), using only the figure-mentioning paragraphs—no images at all. The text-only approach worked remarkably well. We also found that roughly 75% of caption words could be traced directly to those paragraphs or the figure’s OCR.
This work resulted in our INLG 2023 paper, which received the Best Long Paper award and was selected as one of the three finalists for the Best Evaluation Paper award (Huang et al. 2023). It demonstrated that scientific figure captioning behaves much more like contextual summarization than visual description. Furthermore, a small human evaluation in the paper confirmed our motivating observation: many author-written captions were mediocre at best. PhD students rated over 50% of author-written captions as unhelpful. This result highlights the limitations of caption authoring practices and the potential for LLM assistance to improve caption utility for readers.
@@
Against this backdrop, as well as with ChatGPT reshaping the field, we began experimenting with LLMs as automatic judges in early 2023. PhD students and trained undergraduates rated captions generated by different approaches, and we compared their ratings to those produced by LLMs. The results, published at EMNLP 2023 (Hsu et al. 2023), were interesting: LLM judgments aligned well with PhD students’ ratings. Undergraduate ratings, however, diverged. Less-experienced readers cared more about clear take-away messages, while experts prioritized numerical precision and technical detail.
Our study landed at the early wave of the “LLM-as-judge” trend and drew more attention than we expected. Interestingly, a later 2024 study would report that LLM judges sometimes favor their own generations (Panickssery et al. 2025). Our work avoided that bias by deliberately excluding LLM-generated captions from the evaluation pool.
@@
We launched the first SciCap Challenge in 2023 (Figure 1) at the ICCV workshop Closing the Loop Between Vision and Language. We invited teams around the world to generate captions for an expanded version of our dataset, i.e., the SciCap Challenge dataset, which contained 476,389 single-panel scientific figures across 8 scholarly domains and 5 figure types. [^1] The SciCap Challenge dataset (Hsu et al. 2025) included figures from arXiv papers across computer science, physics, math, economics, statistics, quantitative biology, quantitative finance, and electrical engineering. Figures were categorized into node diagrams, equations, graph plots, scatterplots, and bar charts. We limited the per-paper sampling to encourage participation from teams with modest computational budgets. This dataset was more than three times the size of our original 2021 SciCap release. Each figure came with its image, OCR text, figure type, and the paragraphs in the arXiv that explicitly mentioned it. When we designed the specification of the SciCap Challenge, we intentionally made the figure-mentioning paragraphs part of the standard input that every team received. We wanted this design choice to send a clear signal: scientific figure captioning is a highly context-dependent generation task; a good caption cannot be written from an image alone.
Built on the outputs submitted to the 2023 Challenge, we conducted a large-scale human and automatic evaluation study, later published at TACL (Hsu et al. 2025) and presented at NAACL 2025. In this paper, we found that multimodal LLMs (MLLMs), particularly GPT-4V, consistently produced captions preferred by professional editors over all other approaches, including human-written captions. This was true across domains, across figure types, and even for papers published after the models’ claimed training cut-off date. In cases where authors wrote captions carelessly, MLLM-generated captions were often clearly better—a humbling result, but also an exciting one for the potential of assistive tools.
Since 2023, we have continued to run the SciCap Challenge annually—at IJCAI 2024 and at the LM4Sci Workshop at COLM 2025. Over time, this effort has helped form a small yet active community working on scientific figure captioning. Some teams have even returned to participate again after their first participation; we even collaborated with a team to publish their approach (Kim et al. 2025). Our very first SciCap paper (Hsu et al. 2021), which originally was a Findings paper without fancy modeling, has been cited over 130 times.
@@
Each year we ran the SciCap Challenge, we faced the same dilemma: automatic evaluation is cheap and fast but unreliable, while human evaluation is reliable but extremely slow and expensive. Across the 2023-2025 challenges, we experimented with different ways to balance this trade-off. In 2023, we gave two awards: one based on ROUGE-2-Normalized scores on the full hidden test set, and another on a curated subset of higher-quality captions, motivated by our finding that models performing well overall could still fail on well-written captions (Huang et al. 2023). However, rankings barely changed across the two sets. In 2024, we switched to human evaluation on a small sample of 200 figures and split the challenge into two tracks: a short-caption track, where systems had to produce captions no longer than the originals for at least 30% of samples, and a long-caption track with the opposite constraint. This time, the results diverged: teams outperformed the original authors of captions in the long-caption track, but authors still performed better in the short-caption track. This could be due to humans’ known preference for longer captions. In 2025, to mitigate this bias, we went with two distinct tracks, one judged by humans and one judged automatically, using the new LaMP-Cap dataset (Ng et al. 2025b) focused on personalized captioning. This time, the results diverged again across two sets. Teams were able to mimic the author’s style per automatic evaluations, but this often hurt caption quality perceived by humans. In fact, human evaluators consistently preferred GPT-4o’s vanilla captions over personalized ones.
@@
Our first step was building SciCapenter, an interactive caption-writing assistant that displays the figure, its figure-mentioning paragraphs, and several AI-generated draft captions, and allows writers to iteratively revise them with quality ratings and contextual guidance (Hsu et al. 2024). A pilot study suggested that these suggestions reduced cognitive load and helped writers produce higher-quality text. We then conducted a separate, deeper study in which 18 researchers rewrote captions from their own recently published papers while using multiple LLM-generated captions as starting points (Ng et al. 2025a). Through video-based interaction analysis, we found that writers typically began by copying or adapting AI-generated captions, preferred longer and detail-rich drafts, and relied heavily on the textual–visual alignment of the suggestions. However, they struggled when figures were conceptually complex or when AI suggestions conflicted with disciplinary norms. We also saw broader themes (Ng et al. 2025c): caption writing is deeply multimodal, writers often lack linguistic confidence but trust their domain expertise, they benefit from multiple diverse AI suggestions, and caption norms vary dramatically across fields.
@@
In our caption-writing studies (Ng et al. 2025a, c), we noticed a recurring gap: even when AI-generated captions were of fine quality, authors often felt they were too generic and did not match their writing style or the stylistic conventions of their target venues. This pushed us to think seriously about personalization. Although personalized text generation is now common with LLMs (Zhang et al. 2024), the key practical question for our task was: what information can we use to characterize the style of a group of co-authors? Using all papers written by each co-author was an obvious answer, but it was expensive and conceptually messy, since scholarly writing is collaborative and we rarely know which author actually wrote which caption. So we turned to something more grounded and reliable: information within the same paper.
This led us to create LaMP-Cap (Ng et al. 2025b), a dataset derived from the SciCap Challenge dataset and designed specifically for personalized scientific figure captioning using multimodal profiles. For each target figure, LAMP-CAP provides the figure image and its figure-mentioning paragraphs as model input, along with up to three additional figures from the same paper (each with its own image, caption, and paragraphs) as a style profile. Across 110,828 target figures, we tested four LLMs and consistently found that adding these multimodal profiles helped models generate captions closer to the original author-written ones. Human evaluation further confirmed the usefulness of personalization. Interestingly, human judges preferred captions generated with one profile figure over both no-profile and all-profile conditions. This echoes a common dilemma in personalization (Zhang et al. 2024): a little tailoring improves quality and feels helpful to users, but pushing too hard to mimic the user’s style can hurt the generation quality.
@@
Real-world writing rarely unfolds in neat, fully drafted documents. To help authors, captioning systems must handle partial or missing context for figures in unfinished papers or for standalone visuals circulating on social media, in news, or in advertisements. 18.81% of figures in the original SciCap dataset lacked any identified mentions (Huang et al. 2023). Generating meaningful captions when only the image is available remains one of the field’s most practical yet unsolved problems.
@@
SciCap’s evaluations showed that reader expertise shapes caption preference (Hsu et al. 2023). Undergraduates valued clear takeaways; PhD students cared about methodological precision and numbers. Future models might adapt captions for different audiences, such as experts, students, and practitioners, which will require advances in audience modeling, controllable generation, and implicit preference inference. Building audience-aware captioning systems without relying on extensive user profiling remains an open challenge in personalization.
@@
The experimental results from both the LaMP-CAP dataset (Ng et al. 2025b) and the SciCap Challenge 2025 showed a clear trade-off: strong stylistic imitation can reduce factual accuracy or clarity. This tension between stylistic faithfulness and communicative quality raises important questions about authorship, identity, and the design of human-AI co-writing systems. It also brings up a broader question that some NLP researchers have asked: to support writing, do we actually need personalized text generation, or is generic generation sufficient for most authors to build on?
@@
Most caption research evaluates outputs from the reader’s perspective, not the writer’s. Yet to understand how AI truly helps scientists, we must study caption writing in realistic conditions. Recruiting authors actively drafting their own papers is difficult; proxy studies, e.g., rewriting old captions or captioning others’ figures (Ng et al. 2025a, c), only approximate the real task. The Sloan-funded deployment of SciCapenter opens a rare opportunity to observe genuine writing-in-progress (Huang and Fetzer 2024), but conducting such ecologically valid studies remains expensive and methodologically hard.